OpenStudy (anonymous):

verify the identity: cos3x=4cos^3 x-3cosx

OpenStudy (anonymous):

\[\cos3x=4\cos ^{3}x-3cosx\]

OpenStudy (campbell_st):

well start with cos(2x + x) = cos(2x)cos(x) - sin(2x)sin(x) use an appropriate substitution for cos(2x) ans sin(2x) = 2sin(x)cos(x) try cos(2x) = 2cos^2(x) - 1 sp you have \[ ( 2\cos^2 (x) -1)\cos(x) - (2\sin(x)\cos(x))\sin(x)\] which simplifies to \[2\cos^3x - \cos(x) - 2\sin^2(x)\cos(x)\] now use the substitution sin^2(x) = 1 - cos^2(x) so you get \[2\cos^3(x) - \cos(x) - 2(1 - \cos^2(x))\cos(x)\] just distribute the term on the right and collect like terms for the answer.